### v1

# RU-AI: A Large Multimodal Dataset for Machine Generated Content Detection

Abstract.
The recent advancements in generative AI models, which can create realistic and human-like content, are significantly transforming how people communicate, create, and work. While the appropriate use of generative AI models can benefit the society, their misuse poses significant threats to data reliability and authentication. However, due to a lack of aligned multimodal datasets, effective and robust methods for detecting machine-generated content are still in the early stages of development. In this paper, we introduce RU-AI, a new large-scale multimodal dataset designed for the robust and efficient detection of machine-generated content in text, image, and voice. Our dataset is constructed from three large publicly available datasets: Flickr8K, COCO, and Places205, by combining the original datasets and their corresponding machine-generated pairs. Additionally, experimental results show that our proposed unified model, which incorporates a multimodal embedding module with a multilayer perceptron network, can effectively determine the origin of the data (i.e., original data samples or machine-generated ones) from RU-AI. However, future work is still required to address the remaining challenges posed by RU-AI. The source code and dataset are available at https://github.com/ZhihaoZhang97/RU-AI.
multimodal datasets, generative models, content detection
ccs: Information systems Multimedia content creation
1. Introduction
The recent advent of large generative AI models has significantly impacted various sectors of human society, including industry (Cardon et al., 2023; Dwivedi et al., 2023; Epstein et al., 2023), research (Walters and Murcko, 2020; Megahed et al., 2023; Peres et al., 2023), and education (Qadir, 2023; Peres et al., 2023). Current generative AI models are capable of producing realistic images (e.g. Stable Diffusion (Rombach et al., 2022), MidjourneyV5 (Holz et al., [n. d.]), DALL-E2 (Ramesh et al., 2022)), human-level text (e.g. ChatGPT (Ouyang et al., 2022), GPT-4 (OpenAI, 2023), LLaMa (Touvron et al., 2023a)) and natural-sound voice (e.g. YourTTS (Casanova et al., 2022), VITS (Kim et al., 2021)), which covers the major multimedia information in communication, creation and work. Deploying generative AI models appropriately can save professionals’ time for substantive matters (Hacker et al., 2023), assist scientists in discovering new medicine or compounds (Zhavoronkov et al., 2019), and provide students with customised explanations and feedback (Grassini, 2023). However, generative AI models’ capabilities may be misused for generating discriminate or disrespectful content (Borgesius, 2020), spreading misinformation (Zhou et al., 2023), and conducting target scams (Wang et al., 2020a). Hence, it is important to identify whether a piece of content is from a human or a machine.
Figure 1. Data collection process for all datasets. It shows the text collection (left), the image collection (middle), and the voice collection (right). Text is generated through rephrasing caption text. Images are created by referencing both caption text and real images. Voices are synthesized using caption text. All collected data is then matched with their original data pairs.
Previous studies have explored manual methods and automated approaches for detecting machine-generated content. Researchers (Mori et al., 2019; Samo and Highhouse, 2023; AlBadawy et al., 2019) have found that humans can identify machine-generated content with substantial accuracy, but the quality of results is largely influenced by an individual’s background, experience and preference. Moreover, human experts may experience difficulties in distinguishing machine-generated content from real content in specific domains, such as creative writing (Gunser et al., 2021), art painting (Elgammal et al., 2017) and unfamiliar voices (Mai et al., 2023). In addition, identifying machine-generated content manually is labour-intensive and challenging to scale, which results in machine-generated content not being identified promptly. Therefore, it is important to explore automated approaches capable of detecting machine-generated content across different modalities, with a balance between accuracy and efficiency. In practice, corresponding comprehensive datasets form the foundation for developing automated detection approaches and models. However, the majority of the existing datasets for machine-generated content detection (Fagni et al., 2021; Wang et al., 2023; Liu et al., 2023; Li et al., 2023b; Guo et al., 2023; Wang et al., 2020a; Frank and Schönherr, 2021; Salvi et al., 2023; Göring et al., 2023; Wang et al., 2020b; Zhu et al., 2023) are constructed within a single modality, disregarding valuable information across modalities that could aid in detecting the machine-generated content. Although there are several multimodal datasets for machine-generated content detection (Khalid et al., 2021; Dolhansky et al., 2019; He et al., 2021), they mainly contain information on human faces and voices, significantly limiting the scope of their usage. Only one dataset (Shao et al., 2023) contains machine-generated images and text in the general domain, but it still lacks of machine synthesised voices.
To address these problems, we introduce RU-AI, a large-scale multimodal dataset for robust detection of machine-generated content in text, image and voice, and develop baseline detection models and perform extensive experiments with them. The real portion of our multimodal dataset includes three large publicly available image-caption datasets: COCO (Lin et al., 2014), Flickr8K (Young et al., 2014) and Places205 (Zhou et al., 2014), along with their corresponding human voice transcriptions. Machine-generated data is produced by inferencing generative AI models, including five models for text generation, image generation, and voice generation. This yields 245,895 aligned instances each for both real and machine-generated data. Moreover, to prove the effectiveness of our dataset, we develop and train baseline detection models on RU-AI. Our contribution can be summarised as follows:
  • We introduce a newly constructed large-scale multimodal dataset that comprises true data and machine-generated data, with instances aligned across text, image, and voice modalities. The dataset contains multiple categories of objects and the machine-generated data is produced by using five generative AI models for each modality.
  • We develop baseline detection models and conduct experiments by testing the performance of them across all modalities, demonstrating the effectiveness and values of our constructed datasets in training and testing models for machine-generated content detection.
2. RU-AI Dataset
We introduce RU-AI, a new large-scale dataset that features multimodal real and human-generated content, accompanied by machine-generated content. It encompasses three modalities: real-world images, image-describing caption text, and spoken voices for captions, providing more than 245k real and machine-generated aligned instance pairs for each modality. More specifically, image and text instances are collected from COCO (Lin et al., 2014), Flickr8K (Young et al., 2014) and Places205 (Zhou et al., 2014), while voice instances are sourced from SLS Corpora (Harwath et al., 2016; Harwath and Glass, 2017; Hsu et al., 2021; Harwath and Glass, 2015), which form the real and human-generated part of our dataset. The machine-generated portion is produced by five state-of-the-art generative models specific to each modality. Figure 1 presents the data construction pipeline.
2.1. Data Construction
Text. Since COCO and Flickr8K incorporate five captions for each image, we extract the longest caption for each image to ensure the most comprehensive description. For Places205, which includes a single caption for each image, we sort the images’ captions in descending order and extract the first 12,000 image-caption pairs. It is worth noting that Places205 has a large number of captions without punctuation, while only a small number of non-punctuated captions present in Flickr8K and COCO. To maintain consistency across datasets, we utilise FullStop (Guhr et al., 2021) to add punctuation for all non-punctuated captions. The text data is then equally distributed among five Large Language Models (LLMs) for rephrasing. We select both open-source and proprietary LLMs based on different architectures. The rephrased generations are stored with the original captions in a structured data format identified by unique IDs.
Image. To ensure the quality and consistency of machine-generated images across all three datasets, we exclude images under 224 pixels in width or height and resize those over 640 pixels to a maximum of 640 pixels while maintaining their original aspect ratios. We apply five diffusion models with different parameter sizes and fine-tuning techniques to generate images, each contributing equally to the overall dataset. Image captions, positive-negative prompt pairs, and original images are used as seeds for these models during inference to ensure that the semantic meaning and realistic scenes are captured in image generation. The generated images maintain the same aspect ratio as the original images and are named after the original image IDs with the model names attached.
Voice. The voices for captions are cloud-sourced from a group of individuals, with each individual being assigned random captions (Harwath et al., 2016; Harwath and Glass, 2017; Hsu et al., 2021; Harwath and Glass, 2015). Some text-to-speech (TTS) models require reference voices for inference; therefore, we merge the spoken caption voices of each individual, identified by their unique user ID, into one-minute long voice files. For YourTTS (Casanova et al., 2022), the machine-generated voices are synthesised by using the captions along with their corresponding one-minute human voices as references. For other models, the voices are synthesised directly from the caption text using default settings. All five models contribute equally to the generated voice data, and these voices are named and saved similarly to the images, with the original IDs and model names attached.
2.2. Data Processing
LLMs may add extra information such as ”Sure, here is the sentence:” and ”Certainly, here’s a rephrased version of your text:” during the rephrasing process, which is unrelated to the generated sentences. We utilize a regular expression method to remove this information from the generated sentences to ensure the quality of the text data. Since censorship can be applied to generative models for safety and alignment, empty sentences or blank images can appear during the generation process. We remove them to ensure consistency across our dataset. To ensure alignment across the three modalities, we perform a data-matching process shown in Algorithm 1, where real and human-generated content is represented by $R$, and machine-generated content is represented by $N$. The generated images and captions, represented by $N_{image}$ and $N_{caption}$, are a subset of the real images and captions, while $R_{voice}$ corresponds to human-spoken voices, having the same amount of data as the human-generated ones. For each instance, we retrieve the unique IDs of the real images and captions ($R_{image_{i}}$ and $R_{caption_{i}}$) and verify whether they both exist in the generated sets. Since we intentionally ensure that the original and machine-generated data share the same unique ID, a matched data instance is formed by aligning the image, text, and audio that correspond to the specific ID. We implement this approach in our collected dataset, resulting in 245,895 aligned real-and-AI data pairs. As shown in Table 1, the final processed data for each modality includes 7,890 pairs from Flickr8K, 119,264 pairs from COCO, and 118,741 pairs from Places205.
Algorithm 1 Data matching across text, image and voice
$R_{caption},R_{image},R_{voice}$
$N_{image},N_{caption},N_{voice}$
while $N_{image}$, $N_{caption}$ do
     if $N_{image_{i}}\in R_{image}\And N_{caption}\in R_{caption}$  then
         $M_{match_{j}}\leftarrow(N_{image_{i}},R_{image})\And$
         $M_{match_{j}}\leftarrow(N_{caption_{i}},R_{caption_{i}})\And$
         $M_{match_{j}}\leftarrow(N_{voice_{i}},R_{voice_{i}})$
     end if
end while
return $M_{match}$
Table 1. The number of machine-generated data from different datasets in each modality of RU-AI.
Modality Model Dataset
Flickr8K COCO Place
Text ChatGPT (Ouyang et al., 2022) 1,610 24,272 23,736
Llama-2-13B (Touvron et al., 2023b) 1,518 23,121 23,488
RWKV-5-7B (Li et al., 2024) 1,628 23,898 23,780
Mixtral-8x7B (Jiang et al., 2024) 1,599 23,752 23,858
PaLM-2 (Anil et al., 2023) 1,535 24,221 23,879
Image StableDiffusion-1.5 (Rombach et al., 2022) 1,610 24,272 23,736
StableDiffusion-1.5-Hyper (Ren et al., 2024) 1,518 23,121 23,488
AbsoluteReality-1.8  [^1] Both authors contributed equally to this research. 1,628 23,898 23,780
epiCRealism-VAE  [^2] https://huggingface.co/emilianJR/epiCRealism 1,599 23,752 23,858
StableDiffusion-XL (Podell et al., 2023) 1,535 24,221 23,879
Audio YourTTS (Casanova et al., 2022) 1,610 24,272 23,736
XTTS-2(Golge and Team, 2021) 1,518 23,121 23,488
EfficientSpeech (Atienza, 2023) 1,628 23,898 23,780
VITS (Kim et al., 2021) 1,599 23,752 23,858
StyleTTS-2 (Li et al., 2023a) 1,535 24,221 23,879
Total / modality 7,890 119,264 118,741
11footnotetext: https://huggingface.co/Lykon/AbsoluteReality
2.3. Data Characteristics
Table 2. Statistics of generated text data. ”Avg. Len.” is the average number of words per sentence, ”Avg. RL” is the average ROUGE-L score, and ”Avg. BS” is the average BERT Score. Reference is the original text.
Text Source Avg. Len. Avg. RL Avg. BS
ChatGPT 22.97 0.3698 0.9167
Llama-2-13B 26.96 0.2795 0.8852
RWKV-5-7B 21.07 0.6411 0.9464
Mixtral-8x7B 19.28 0.5662 0.9411
PaLM-2 17.23 0.4615 0.9278
Reference 21.66 N/A N/A
Table 3. Statistics of synthesised voice data. ”Avg. Len.” is the average duration of the voices in seconds, ”Avg. AR” is the average amplitude ranges, and ”Avg. BS” is the average spectral flatness. Reference is the original voices.
Voice Source Avg. Len. Avg. AR Avg. SF
YourTTS 7.338 1.830 0.1264
XTTS-2 7.842 1.826 0.1140
EfficientSpeech 5.980 1.416 0.02997
VITS 7.536 1.831 0.09121
StyleTTS-2 7.563 1.001 0.03474
Reference 8.700 1.423 0.03605
Our dataset includes 245,895 pairs of real and machine-generated data from text, image, and voice modalities, resulting in a total of 1,475,370 instances collected from three diverse datasets. Table 2 shows the average length of captions from different LLMs and their average similarity between the reference and generated text, indicated by average ROUGE-L and BERTScore. Table 3 shows the synthesised voice characteristics from different models, including average voice duration, average amplitude, and average spectral flatness. The average amplitude range refers to the average normalised audio scale, while the average spectral flatness is a measure of the sound’s noise-like level. Figure 2 displays the object categories featured in images, which also serve as the topics in text and voice modalities. Since all images in Flickr8K are unlabelled with object instances, we employ pre-trained YOLOv5 (et. al., 2021) to annotate the unlabelled images. The data characteristics are based on the original annotations and the YOLOv5 annotations. Our dataset can be divided into 20 different categories, with ”Person” being the largest category, and there are 1,040 images that do not belong to any category and are classified as ”Other”.
Figure 2. Distribution of image data in each category.
3. Proposed Baseline
Based on the RU-AI dataset, we propose a unified classification model capable of determining the origin of input data from different modalities. Our model utilises a multimodal embedding model to project the features from different modalities including text, image and voice, into a common space through high-dimensional embedding sequences. These embedding sequences are then fed into a multilayer perceptron network to distinguish the origin of the input data.
Implementation Details. We use two pre-trained multimodal encoder models with their large settings as the baseline multimodal embedding models: ImageBind (Girdhar et al., 2023) and LanguageBind (Zhu et al., 2024). ImageBind encodes all input data into 1,024-dimensional embedding sequences, while LanguageBind encodes them into 768-dimensional embedding sequences. Each multimodal embedding model is connected to a separate multilayer perceptron network, consisting of a single 256-dimensional hidden layer. The input layers of these networks match the dimensions of their respective embedding models, 1,024-dimensional for ImageBind and 768-dimensional for LanguageBind. These networks are used for the classification of the input data as either real or machine-generated one. Specifically, the output is calculated from probability vectors $\vec{z}=(z_{r},z_{m})$ using the softmax function $\sigma(\vec{z})_{i}$. The instance is classified as real if $z_{r}>z_{m}$; otherwise, it is classified as machine-generated. To analyse the effectiveness of different modalities in our data, we train and evaluate our classification model with different combinations of the data. The training samples are randomly selected from 80% of each data combination, while the remaining 20% of the samples are used as testing sets. We maintain the same size of data throughout the experiment and distribute the modalities equally. The performance is evaluated using Accuracy, Precision, and F1 metrics on the trained-modality (TM) and all-modalities (ALL). For the baseline implementation, we use PyTorch 1.13 with CUDA 11.8 on a server with a single Nvidia RTX 3090 GPU. All models were trained for 5 epochs, using the Adam optimizer with cross-entropy loss and a learning rate of 1e-4, with the embedding model frozen. We follow the embedding models’ original feature extraction techniques in all three modalities for consistency across datasets.
Results.The classification results from our proposed model are shown in Table 4. We observe that voice is the simplest modality to distinguish between human and machine-generated content. This aligns with the model behaviours in all-modality evaluation, as the accuracy and precision are the lowest when the model is trained only with voice data. This is likely attributable to the obvious feature difference between human and machine-generated voices. Combining data from different modalities improves the model’s performance in all-modality evaluation but affects its effectiveness on the trained modality. This is expected since mixing modality data is more complex than single modality data, which could cause difficulty in model convergence. Models utilising LanguageBind as embedding models generally perform better than those incorporating ImageBind as embedding models. This might be due to LanguageBind’s enhanced cross-modality architecture, which binds all modalities through language with rich semantics to provide more comprehensive information even with a smaller embedding size. According to our baseline results, machine-generated content detection remains a challenging task, especially in terms of text and cross-modal detection. To subside the potential misuse of generative models, future work is required to explore more robust and reliable methods for detecting machine-generated content.
Table 4. Results of the proposed classification model. ”LB” is LanguageBind, ”IB” is ImageBind, ”A” is Accuracy, ”P” is Precision, ”ALL” is all three modalities, ”TM” is trained-modality.
Model Modality A@TM P@TM F1@TM A@ALL P@ALL F1@ALL
LB + MLP Text 75.04 75.02 75.04 69.76 69.78 69.75
Image 94.74 93.78 94.79 76.49 85.95 72.93
Voice 100.00 100.00 100.00 68.89 64.09 73.43
Text+Image 74.05 73.15 74.55 77.09 76.91 77.16
Text+Voice 86.25 85.57 86.38 77.69 81.56 76.23
Image+Voice 96.97 96.88 96.97 82.26 78.89 83.23
All N/A N/A N/A 84.20 84.20 82.67
IB + MLP Text 80.83 79.91 81.12 60.70 57.36 67.96
Image 95.23 94.66 95.26 67.75 71.03 65.03
Voice 100.00 100.00 100.00 65.77 63.61 68.28
Text+Image 79.37 77.97 79.87 71.48 75.18 69.23
Text+Voice 87.95 87.97 87.95 75.39 69.05 78.90
Image+Voice 96.97 97.02 96.96 81.25 82.84 80.78
All N/A N/A N/A 80.63 79.19 81.10
4. Conclusion
We propose RU-AI as a large multimodal dataset for machine-generated content detection. Compared with other datasets that contain only single or dual modalities, our approach aligns text, image, and voice data with their machine-generated counterparts. To ensure data diversity, we incorporate three datasets across 18 different object categories, using five generative models for each modality. Experimental results from our proposed baseline show that our unified classification model can effectively detect machine-generated content. However, several challenges remain, such as improving cross-modality detection accuracy, addressing model scalability and availability to general public. We hope making RU-AI dataset publicly available to the public provides direction for future research.
5. Ethics Considerations
The collection of RU-AI dataset is in accordance with institutional review board standards and ACM standards of ethics. Our dataset is created from open-sourced datasets and generative models with respect for the privacy and confidentiality of original sources. We publish our dataset at Zenodo with Digital Object Identifier: 10.5281/zenodo.11406538 to encourage open and collaborative research.